a baseball is hit almost strait up a kangroo jumps strait up to a vertical height of 1.75 m. how long was it in the air before returning to earth
Answer #1

The initial velocity of the Kangaroo from equation of motion is, v2 =u? –2g(v-yo) 0=u² -2g(y- y.) u = 2g(y – yo) u= 128(y – y